OpenStudy (anonymous):

A husky dog does 20 kJ of work as he pulls a sleight with a force of 340 N. Through what distance does he pull it?

OpenStudy (michele_laino):

hint: the magnitude \(W\) of work done by a force \(F\), is given by the subsequent formula: \[\huge W = F \times s\] where \(s\) is the displacement of such force, and providing that, both force and displacement have the same direction. So we have: \[\huge s = \frac{W}{F} = \frac{{20000}}{{340}} = ...meters\]

OpenStudy (michele_laino):

more precisely, \(s\) is the magnitude of the displacement